### v1

# Search-Based Software Engineering in the Landscape of AI Foundation Models

Abstract
Search-based software engineering (SBSE), at the intersection of artificial intelligence (AI) and software engineering, has been an active area of research for about 25 years. It has been applied to solve numerous problems across the entire software engineering lifecycle and has demonstrated its versatility in multiple domains. With the recent advancements in AI, particularly the emergence of foundation models (FMs), the evolution of SBSE alongside FMs remains undetermined. In this window of opportunity, we propose a research roadmap that articulates the current landscape of SBSE in relation to foundation models (FMs), highlights open challenges, and outlines potential research directions for advancing SBSE through its interplay with FMs. This roadmap aims to establish a forward-thinking and innovative perspective for the future of SBSE in the era of FMs.
Keywords AI  $\cdot$ Search-based Software Engineering  $\cdot$ Metaheuristic Optimization  $\cdot$ Foundation Models  $\cdot$ Large Language Models  $\cdot$ Vision Language Models  $\cdot$ Multimodal Models
1 Introduction
Search-based software engineering (SBSE) has been a prominent field for nearly a quarter-century—approaching its silver jubilee—since it was first introduced by Harman and Jones [17] in 2001. Over the years, it has rapidly evolved to address emerging and complex software engineering problems. It has been successfully applied throughout the software engineering (SE) lifecycle, including requirements engineering, software design, software development, software testing, deployment, and maintenance [19].
SBSE uses metaheuristic search optimization techniques to solve SE problems. This involves reformulating SE problems as metaheuristic search problems by creating solution representation, defining fitness functions, and selecting search operators [17]. The problem formulated in this way is then solved by applying search and optimization techniques to identify optimal solutions. These techniques include single-objective, multi-objective, and many-objective search algorithms, each designed for different problem types and levels of complexity [18, 19].
Recently, artificial intelligence (AI) foundation models (FMs) have brought a major transformation, attracting widespread interest across research communities, academia, and industry [23]. These models are trained on vast amounts of data and are capable of performing a wide range of analytical tasks. For specific application domains, these models can be adapted using techniques such as fine-tuning and prompt engineering. Depending on their modalities, architecture, and application domains, FMs can be categorized as follows: Large Language Models (LLMs) like GPT for textual content; Vision Language Models (VLMs) like ResNet for image and video data; Speech and Audio Models (SAMs) like WaveNet for tasks related to speech recognition, synthesis, and audio analysis; and Multimodal Models (MMs) like CLIP that integrate multiple types of data, such as text and images, to enable content generation and comprehension across different modalities [44]. Their ability to generalize knowledge across multiple domains and adapt to diverse applications has enabled new AI-driven innovations, making them a focal point of exploration. As a result, researchers are actively exploring their potential, academic institutions are integrating them into curricula, and industries are leveraging them to enhance automation, decision-making, and user experiences.
Considering rapid AI transformations driven by FMs, in this paper, we present a strategic roadmap exploring the synergistic interplay between SBSE and FMs in the context of future innovations and advancements. We structure this roadmap to explore four key dimensions, as illustrated in Figure 1: (i) how FMs can be leveraged for SBSE design (Section 2), (ii) how FMs can complement SBSE applied for SE problems (Section 3), (iii) how well-studied SBSE practices can be adapted to FMs customized for particular SE activities (Section 4), and (iv) how SBSE and FMs operate together and their synergistic potential (Section 5). For each dimension, we review the current state of the art, discuss open challenges, and present opportunities with a forward-looking vision that outlines potential research directions.
Figure 1: Research roadmap dimensions outlining the potential integration of SBSE with AI foundation models.
2 Foundation Models for Automating Search-based Software Engineering Design
Recently, LLMs have attracted significant interest in evolutionary computation due to their ability to enhance various aspects [42]. An important step in this regard is to use LLMs as search operators for single and multi-objective optimization, such as language model crossover (LMX) for single-objective [28] and LLM-assisted offspring generation for multi-objective optimization [41]. Another line of work is evolutionary algorithm generation and improvement using LLMs. Notable works include using LLMs to generate swarm intelligence algorithms [31], designing heuristic algorithms [24], and evolving search algorithms [45]. In addition, some studies also focused on using LLMs to assist novice users in interpreting and comprehending evolutionary algorithm-generated solutions [25, 38]. While LLM-based evolutionary computation has recently gained attention in the literature, its effectiveness within the SBSE context remains understudied. Moreover, given LLMs’ potential to enhance evolutionary computation, we anticipate that other FMs, such as VLMs and MMs, could further complement SBSE in solving complex SE problems. Considering this, we outline key research opportunities as follows:
2.1 Design of Fitness Functions for SBSE Problems
Typically, a software engineer manually identifies and defines the fitness functions for SBSE problems based on their understanding of the domain. FMs have the potential to assist in automating the definition and subsequent implementation of fitness functions for a given SBSE problem. For example, in test case prioritization, FMs can help provide alternative fitness function options to guide search algorithms in prioritizing test cases, e.g., based on historical data on faults found by each test case and the coverage achieved. We foresee the following options to be investigated with FMs for designing fitness functions for SBSE problems: 1) FMs can be studied as a chatbot created for SBSE users, where they can build fitness functions by interacting with the chatbot and then implement their fitness functions themselves; 2) SBSE users can provide a set of requirements and ask an FM to implement the fitness function. In addition, VLMs, SAMs, and MMs can be used to derive fitness function suggestions from images, audio, and video data, for example, when generating test scenarios for autonomous driving systems (ADS).
2.2 Design of Search Operators
In a recent study, Wu et al. [42] categorized various LLM-based search operators, such as crossover and mutation, for single and multi-objective algorithms. However, the application of such operators in SBSE remains undetermined. Therefore, one potential research direction is to explore LLM-based search operators in the SBSE context. Another research direction is to define search operators using other FMs like VLMs and MMs to enable SBSE applications to emerging domains such as autonomous systems and quantum computing. For example, when dealing with ADS data in image format, VLM-based search operators can be employed to generate test scenarios. Similarly, for ADS data in audio or video formats, search operators based on SAMs and MMs can enhance test scenarios considering the mixed-mode nature of data.
2.3 Solution Encoding for SBSE Problems
A fundamental step in SBSE is solution encoding for a particular SE problem. This varies from one problem nature to another and requires problem-domain knowledge. FMs with diverse capabilities can support solution encoding for different SE lifecycle phases. For instance, to verify software models with SBSE during the design phase, MMs with both textual and visual capabilities can facilitate solution encoding by analyzing software requirements and design artifacts, often created using graphical modeling languages such as Unified Modeling Language (UML).
2.4 FMs for SBSE Implementation Generation
One advanced application of FMs is generating either partial or complete implementations of SBSE problems. In a partial implementation, FMs generate code to solve an SBSE problem, which the user then completes manually. Alternatively, FMs can create complete implementations of SBSE problems that can be further refined. This capability extends across key SBSE components, including solution encoding, operator selection, fitness function design, and search algorithm selection, enabling more efficient and automated problem-solving.
2.5 FMs for Testing, Debugging, and Repair of SBSE Implementation
The implementation of SBSE problems is challenging to test due to the absence of test oracles. To this end, FMs can be used to automate the testing of SBSE problems, ensuring that their implementation is correct. Therefore, further research is needed to assess whether FMs can be used to generate tests as well as test oracles for a given SBSE problem. Once issues are identified in SBSE implementations through testing, it is important to debug and repair them. These tasks are typically performed manually. In this context, FMs can be used to automatically debug SBSE implementations and generate patches to fix identified issues. However, this requires further research investigation, which has not been a focus in the literature.
3 Foundation Models for SE Artifacts Generated with Search-based Software Engineering
In recent years, there has been growing research interest in various SE areas in using LLMs for SBSE. One notable area is software genetic improvement (GI). In this context, Kang and Yoo [20] presented the initial concept of generating mutants from LLMs to optimize the process of GI. Similar to this work, Brownlee et al. [5] employed LLMs as mutation operators to increase diversity in the search process for GI. Software program refactoring represents another emerging direction where FMs are applied in SBSE. In this regard, Choi et al. [10] utilized LLMs to guide a search-based approach, enhancing refactoring efficiency while reducing complexity. Another key area of interest is software testing, where LLMs are being explored for search-based software testing (SBST). In this case, Turhan [40] introduced evoLLve’M, an approach that utilizes LLMs to enhance JUnit test assertions generated by EvoSuite’s evolutionary process. Similar to this work, Biagiola et al. [4] employed LLMs to improve the readability of unit tests generated by EvoSuite through search techniques.
In summary, the literature review indicates that (i) LLMs, among various FMs, are increasingly used for SBSE, and (ii) numerous SE phases, including requirements engineering, software design, and debugging, remain largely unexplored. Despite SBSE’s long-standing role in solving SE problems, several challenges persist, which we anticipate can be addressed through FMs in the future. Below, we highlight key opportunities and research directions.
3.1 FMs for SE Artifact Interpretation
An SE problem in SBSE is formulated using different solution encoding schemes, such as Binary or Tree-based encoding, depending on the problem’s nature. The final output is generated as an encoded solution, often requiring manual effort for interpretation and transformation into a specific SE artifact. This effort varies based on the SE lifecycle phase. For instance, in the case of requirements prioritization for the next release problem (NRP), the SBSE-generated prioritization sequence needs to be interpreted in textural requirements and utilized for NRP. Similarly, software models generated with SBSE need to be converted to graphical models. In this context, FMs can play a significant role in two directions. First, LLMs with text comprehension and analytical capabilities can be used for SBSE-generated SE artifacts that are required to be interpreted and represented in a textual format. Second, VLMs and MMs can be employed to analyze and transform SBSE-generated SE artifacts into graphical representations.
3.2 FMs for SE Artifact Refinement
For some SE problems, SBSE-generated solutions do not need to be transformed. For example, test data generation for primitive types and program improvement with generic programming. Such scenarios require domain experts to verify the generated solution manually. In the case of test data generation, test effectiveness using metrics like coverage needs to be analyzed after running tests on the system under test. Similarly, programs generated with generic programming either need to be compiled and tested, or they need to be manually examined using code inspection and walkthrough techniques. FMs have significant potential in these cases. For the test data generated using SBSE, LLMs can conduct a static analysis utilizing the source code and test data to provide insights on test effectiveness, such as untested paths or incomplete coverage scenarios. In developing LLM-based static analysis techniques, we foresee the potential for devising novel test effectiveness metrics that complement these techniques. Furthermore, LLM-based static analysis (e.g., type checking and data flow analysis) can be further extended to verify programs generated with SBSE statically.
3.3 FMs for Pareto Analysis
In the case of multi- and many-objective SE problems, SBSE generates non-dominated or Pareto-optimal solutions in the objective space. A common practice is to analyze the resultant solutions and select the best one meeting a particular set of requirements. This process needs to be repeated whenever requirements change. In this regard, LLMs can support domain experts in understanding the trade-offs between different objectives and selecting the best solution for a given scenario. Moreover, for non-expert users, VLMs or MMs can construct a visual representation of the Pareto front, illustrating the set of all non-dominated solutions and offering suggestions to assist in selecting the most suitable option.
4 Search-based Software Engineering for SE Artifacts Generated with Foundation Models
FMs are increasingly being integrated into the entire SE lifecycle [23]. This results in FMs customized for specialized SE activities. A prominent example is LLMs tailored for different coding tasks, such as CodeX [9] for code generation and CodeGen [9] for program synthesis. In addition to code generation, these LLMs are employed to generate test cases [13]. For debugging, LLMs have been utilized for fault localization [43] and software bug reproduction from bug reports using prompting techniques [21]. Another line of research is focused on applying SBSE techniques to optimize FMs for specific SE problems. A noteworthy area is software effort estimation, where Tawosi et al. [39] introduced the idea of applying search techniques to optimize few-shot methods for fine-tuning LLMs to enhance their performance in estimating software effort. Another area of research focuses on enhancing image quality through SBSE. In this regard, Berger et al. [3] presented StableYolo, which employs a multi-objective search to optimize LLM prompts and parameters for generating high-quality and realistic images using tools such as YOLO and Stable Diffusion. Later, Gong et al. [16] introduced GreenStableYolo, which applies a multi-objective search for efficient LLM inference while maintaining high-quality image generation.
A key limitation noted in the literature is the lack of sufficiently large, high-quality datasets. Since FMs like LLMs are trained on enormous amounts of data, applying SBSE to optimize techniques such as prompting or few-shot learning necessitates access to similarly extensive and reliable data. Despite this challenge, an overview of existing works on applying SBSE to LLMs indicates that SBSE has significant potential for optimizing SE tasks utilizing FMs. In summary, this highlights several open problems that need to be addressed and significant opportunities for future research to expand the application of SBSE techniques to a broader range of FMs (e.g., VLMs and MMs) and diverse SE domains. Below, we present possible opportunities and future directions.
4.1 Search-based Code Improvement
Recent literature has shown significant interest in using general and fine-tuned LLMs for code generation. However, LLM-generated code often leads to compiler errors, and testing such code remains an open challenge [13]. Furthermore, integrating LLM-generated code in the software code base requires significant manual effort in analyzing code, resolving conflicts, and testing the integrated code. The current practice to improve the code generated by LLMs is using prompt engineering [13]. Building on the extensive research and practical applications of search-based program improvement and code repair, applying these techniques to refine LLM-generated code opens significant research opportunities. This advancement will extend the applicability of traditional SBSE techniques from human-written code to optimizing LLM-generated code.
4.2 Search-based Testing of FMs-Generated Artifacts
Similar to code generation challenges, the test cases generated from LLMs suffer from several issues, such as erroneous tests, low coverage, and test flakiness [13]. The existing approaches tackle these issues using different LLMs and prompting techniques. Given the well-studied SBST techniques, one potential research direction is to apply SBST to improve tests generated with LLMs. Moreover, in case a large number of tests are generated with LLMs, it is infeasible to run all tests for regression testing in continuous integration/continuous deployment (CI/CD) or DevOps environments. Therefore, another research direction is to apply SBST regression test selection, minimization, and prioritization techniques for LLM-generated tests.
4.3 Search-based Debugging, Repair, and Evolution of FMs-Generated Artifacts
LLM-based fault localization and reproduction techniques generate fault reports, which are subsequently used for tasks such as debugging, program repair, and test evolution. Given that SBSE literature has well-established techniques for software maintenance and evolution, we argue that these techniques can be leveraged to enhance the process. For instance, search-based program repair techniques can be adapted to improve code as a next step after fault localization by LLMs. Moreover, once the code has been improved, another direction involves using SBSE to evolve tests based on the fault information.
5 Interplay between Search-based Software Engineering and Foundation Models
A growing research trend is to explore the synergy between SBSE and LLMs, investigating how LLMs can enhance SBSE techniques and how SBSE can improve LLMs for SE tasks. One notable area is software development, where Gao et al. [15] introduced SBLLM, a framework for code optimization that leverages LLMs to apply evolutionary search techniques like crossover and mutation. For the application-specific use of search and LLMs, Shojaee et al. [37] presented LLM-SR, an approach that combines LLMs and evolutionary search to discover mathematical equations from the programs, demonstrating the potential to enhance the accuracy and efficiency with both LLMs and search techniques. Another prominent area of interest is exploring the potential of LLMs with SBST. In this direction, Dakhama et al. [12] introduced SearchGEM5, an approach that combines search-based fuzzing with LLMs to generate test cases for the Gem5 simulator. Later, Bruce et al. [6] presented an approach that utilizes LLMs to guide search-based fuzzing toward finding faults in ARM simulator software. Lemieux et al. [22] presented a technique called CodaMosa, which integrates a multi-objective search algorithm with a coding LLM to generate tests for Python code capable of achieving high coverage. Furthermore, Sapozhnikov et al. [33] presented TestSpark, a tool that utilizes the search-based capabilities of EvoSuite and LLMs’ textual capabilities to generate Junit tests. Subsequently, Abdullin et al. [1] conducted a comparative study of search-based tools, symbolic execution, and TestSpark configured with various LLMs. Their results demonstrated that while LLM-based tools underperformed in the case of code coverage, they outperformed in achieving higher mutation scores.
While the synergistic potential between SBSE and FMs remains largely unexplored, early research reveals a promising interplay. However, it also highlights a significant performance limitation due to the substantial computational resources required by the integrated use of SBSE and FMs. In the following, we present a research roadmap and opportunities, exploring the potential of SBSE and FMs in advancing metaheuristics search in SBSE, overcoming FM-related challenges, solving SE problems, and addressing challenges in emerging domains.
5.1 Advancing Metaheuristics Search in SBSE
As highlighted in the latest survey [42] and our literature analysis, FMs, specifically LLMs, can support various search algorithms in multiple ways. This includes using FMs to generate population, perform crossover and mutation, act as fitness functions, and support multi-objective optimizations. Given the advantages that FMs bring to the search process, one potential research direction is developing novel FM-inspired search algorithms. Furthermore, search-based algorithms, such as genetic algorithms (GAs), initiate the search process with a randomly generated population. During the search process, different selection strategies like tournament and rank-based selection are used to pick individuals from the current population to generate offspring for the next generation. For solving complex domain problems, intelligently selecting and reusing solutions individuals has demonstrated to be an efficient strategy [35, 36]. In this context, a promising research direction is using FMs to leverage domain knowledge for generating an optimized initial population, thereby enhancing the efficiency of the search process. For instance, for SBSE applied to emerging domains requiring mixed-mode capabilities, such as robotics and ADS, FMs like VLMs or MMs can support generating the initial population. This could significantly broaden SBSE’s applicability to diverse, complex domains with cutting-edge challenges. Moreover, during the search, FMs can be used to calculate fitness, depending on the SBSE problem at hand. For example, in ADS testing, FMs can be queried to assess whether a scenario is realistic. This information can then be used to define fitness functions that guide search algorithms for SBSE problems. Thus, more research is needed to investigate which SBSE problems can utilize FMs to calculate fitness during the search. Another significant challenge is that search algorithms often generate invalid solutions that require repairs. To address this, FMs can repair solutions during the search, ensuring the search is properly guided. To this end, new research methods based on FMs are needed to prompt FMs to repair solutions when issues are encountered during the search.
5.2 Addressing FM Challenges
A key challenge of FMs is non-determinism, where the same input can produce varying outputs, affecting consistency and reliability. This non-determinism leads to several challenges in applying FMs in SE development phases. For instance, in CI/CD and DevOps, variability in generated configurations may lead to inconsistencies in software releases. SBST can complement handling non-determinism in FMs by search-based robustness testing techniques with the objective of exploring inputs leading to high model variability. Another direction in which SBSE can be applied is to find the best configurations for software and hardware resources that minimize variability in model behavior. Another challenge of FMs is their inherent uncertainty. Since FMs are trained on vast amounts of data, fine-tuning them also requires a significant amount of data, and the quality of such data (e.g., redundant features or imbalanced data) can be a potential source of uncertainty. Such uncertainties can significantly affect model interference, leading to ineffective performance for several SE phases. SBSE can be applied to select an optimal set of features, which will result in effectively managing data-related uncertainty and efficient training/fine-tuning of FMs. Furthermore, a key challenge in FMs is hallucination, where the model generates entirely fabricated or factually incorrect information [32], posing a significant concern in FM applications. As a result, detecting and mitigating hallucinations has become a key focus of ongoing research [8]. In this regard, we foresee significant potential for SBSE to effectively address the challenge of hallucinations in FMs. Specifically, SBSE can complement black-box methods for hallucination detection, as highlighted by the need for such methods [8]. By leveraging a search-based approach, a search algorithm can iteratively generate a diverse population of prompts, collect corresponding responses from FMs, and use a fitness function to continuously assess response variations, enabling more precise hallucination detection.
5.3 Solving SE Problems
SBSE has been applied to many software requirement engineering (RE) problems, including requirement selection and prioritization, requirement satisfaction, and fairness analysis in requirements. Software requirements are typically elicited and documented in a textual format. FMs, specifically LLMs with text generation, comprehension, and analysis capabilities, can complement search-based RE in many ways. For instance, in the case of requirements selection, LLMs can be provided with software requirements and then act as a fitness function during search optimization for solving the next release problem. Moreover, for requirements prioritization with multi-objective search, LLMs can be used to analyze conflicting objectives and select Pareto-dominant solutions. In addition to RE, SBSE techniques can be applied to test FMs used to solve SE problems (e.g., code generation and test generation). Taking the example of FMs generating source code, the aim is to test whether FMs generate correct code. For this purpose, SBST can be applied to generate test cases that prompt FMs to generate source code. The generated source code can then be used to compute a fitness function, guiding the search to identify faults that make FMs generate the wrong code. To this end, with carefully designed fitness functions, search-based testing techniques can be used to identify faults in FMs for various software engineering tasks. Once faults are identified using SBSE approaches, they can also be employed to repair FMs. This can be achieved by defining a fitness function to generate patches (e.g., fine-tuning parameters) for FMs until the faults are resolved. In software testing, graphical user interface (GUI) testing presents a compelling application for integrating SBSE and FMs. Modern software applications are designed with advanced GUIs to facilitate user-friendly interactions. GUI testing is essential to ensure functionality, visual consistency, compatibility, and usability. The available tools for GUI testing, like GUITAR [30] and Selenium [7] require test scripts to execute tests. Therefore, combining SBST with MMs, which possess mixed-mode capabilities, such as textual and graphical elements, presents an intriguing direction for future research.
5.4 FM-Integrated SBSE for Emerging Domains
FMs integrated with SBSE can be used to solve complex and dynamic nature problems in emerging domains. For example, in system-level testing of live Internet of Things (IoT) applications, the comprehension and analytical capabilities of FMs (such as LLMs or MMs), combined with SBST, can enable testers to perform online testing in an organized and intelligent manner. Similarly, VLMs and MMs can facilitate the safety analysis of self-adaptive robotics by processing mixed-mode data from cameras, sensors, and actuators. Overall, we envision that integrating FMs with SBSE will offer substantial support across the software engineering lifecycle in diverse domains, including cyber-physical systems, smart grids, and autonomous vehicles.
6 Related Works
To the best of our knowledge, this paper presents the first forward-looking roadmap for SBSE in the context of FMs. While several existing studies have presented the state of research and future directions, many are outdated or no longer reflect the latest advancements, specifically the ones due to FMs. Our work complements the existing works by introducing a timely and forward-thinking vision for the evolution of SBSE in the FM era. Initially, Harman et al. [19] presented a detailed survey of the SBSE research state till 2010, challenges, and future directions. Subsequently, Colanzi et al. [11] presented a systematic mapping study to analyze the evolution of SBSE from 2009 to 2019. Sarro [34] later presented a brief vision highlighting SBSE applications for building responsible software systems. Recently, Wu et al. [42] presented a literature review and a roadmap for integrating LLMs with evolutionary algorithms (EAs). In comparison, our work focuses on the synergistic potential between SBSE, which leverages EAs for solving SE problems, and FMs, which encompass not only LLMs but also VLMs, SAMs, and MMs. Furthermore, we specifically present a roadmap focusing on the potential interplay between SBSE and FMs within the SE context, whereas their work focused on cross-domain research areas.
In the context of SBST, McMinn [26] conducted a comprehensive survey that focused on test data generation using search techniques. Similarly, Ali et al. [2] provided a systematic review of the literature on empirical studies related to search-based test case generation up to the year 2007. McMinn [27] presented a review of the SBST area, including studies till 2011, open challenges, and future directions. Neelofar et al. [29] analyzed the strengths and limitations of automated techniques in SBST, specifically focusing on instance space analysis. In recent work, Fraser and Arcuri [14] outlined a brief vision for the future of search-based test generation in the context of LLMs. In contrast, our work takes a broader perspective, exploring the interplay between SBSE and FMs (in addition to LLMs).
7 Conclusion and Future Work
Our analysis of the current state of the art revealed that most research focuses on applying large language models (LLMs) in evolutionary computation and search-based software engineering (SBSE), leaving many research opportunities to explore other foundation models (FMs) like vision language models and multimodal FMs. We also noted some initial, yet immature, efforts in using LLMs for SBSE, SBSE for LLMs, and their combined application to solve various software engineering problems. Building on these insights, we presented a roadmap outlining forward-looking opportunities and research directions, focusing on the potential synergy between FMs and SBSE. In the future, we plan to extend this roadmap in two directions: (i) enhancing the potential of SBSE and FMs across the entire software engineering lifecycle and (ii) exploring application-specific synergies between SBSE and FMs in emerging domains such as quantum computing, extended reality, and autonomous systems.
Acknowledgments
This research is supported by the RoboSAPIENS project, funded through the European Commission Horizon Europe program under Grant Agreement No. 101133807.